To find the pKa of x-281, you prepare a 0.076M test solution of X-281 at 25.0 degrees Celsius. The pH of he solution is determined to be 3.00. What is the pKa of X-281?
Answer #1

Answer:

Concentration (C) = 0.076 M

Temperature (T)= 250C

PH = 3

PH = 1/2Pka -1/2 logC

3 = 1/2 Pka - 1/2 log [0.076]

3 = 1/2Pka - 1/2 x -1.119

3 = 1/2 Pka - 0.559

1/2 Pka = 3.559

Pka = 3.559 x 2

Pka = 7.118